OpenStudy (anonymous):

2x^2 + 8x = -8 need help

OpenStudy (anonymous):

\[2x^2 + 8x = -8\]

OpenStudy (anonymous):

First bring the -8 to the right hand side. Now divide both sides by 2 Now solve the quadratic equation

OpenStudy (anonymous):

sorry left hand side

OpenStudy (anonymous):

What is the quadratic equation?

OpenStudy (anonymous):

divide by 2 x^2 + 4x = -4 x^2 + 4x + 4 = 0 x^2 + 2x + 2x + 4 =0 x(x+2) + 2(x+2) = 0 (x+2)^2 = 0 x = - 2

OpenStudy (anonymous):

yeah just like @apple_pi showed.

OpenStudy (anonymous):

thanks guys